OpenStudy (anonymous):

The function f(x) = 16(5)^x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (anonymous):

Can someone help me get started?

OpenStudy (anonymous):

If it helps here are the answer choices but I need to know how to do the problem in case I get another like it on my test f(x) = 8(5)^x; growth rate 500% f(x) = 8(5)^(2x), growth rate 8% f(x) = 8(2.24)^x; growth rate 8% f(x) = 8(2.24)^(2x); growth rate 224%

OpenStudy (anonymous):

@surjithayer can you help?

OpenStudy (anonymous):

\[5^x=t ^{2x},\] \[5^x=\left( t^2 \right)^x\] \[t^2=5,t=\sqrt{5}=2.24 \] \[new~ rate=2.24\] i think it can be 4th but why 8 in place of 16,i don't know.

OpenStudy (anonymous):

I think it is the 4th too because 2.24=224% and she wants the formula to calculate two times a year so that would be 2x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw